OpenStudy (anonymous):

Given the exponential function f(x)=(3/2)^x, what would be a vertical shift of -2?

OpenStudy (campbell_st):

this the function -2 so y = (3/2)^x - 2

OpenStudy (campbell_st):

cute the y axis at x = -1

OpenStudy (anonymous):

why wouldn't it be +2?

OpenStudy (campbell_st):

think of it this way (3/2)^0 = 1 y intercept at y = 1 you want to move the graph down 2... so (3/2)^0 - 2 = 1 - 2 y intercept at y = -1 to move up add 2 (3/2)^0 + 2 1 + 2 y intercept at y = 3

OpenStudy (anonymous):

ah..i got it now...my teacher really confused me on this step

OpenStudy (anonymous):

thanks